OpenStudy (anonymous):

Algebra 2 help? Please

OpenStudy (anonymous):

Hi welcome to Openstudy! Can you post your question?

OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)).

OpenStudy (anonymous):

I really need help with this

OpenStudy (anonymous):

it means that the independent variable of f(x) will be the dependent variable of g(x) or in other words input of f(x) = output of g(x). for the x value of f(x), substitute the function g(x) it will be like this: \( f(x) = -4x + 7,\ g(x)=10x - 6\\f(g(x))=-4(10x - 6)+7\) then simplify it

OpenStudy (anonymous):

@bblock5 do you understand?

OpenStudy (anonymous):

dont understand the substitution part

OpenStudy (anonymous):